OpenStudy (anonymous):

I need help with a couple math problems, they all have to do with slope intercept form look below for questions please

OpenStudy (anonymous):

1. x-3y=-18

OpenStudy (anonymous):

Slope intercept form is y=mx+b

OpenStudy (anonymous):

2. 5x +4y = 100

OpenStudy (anonymous):

I know that I need to know what the answer is

OpenStudy (anonymous):

3. 9x = 2y + 13

OpenStudy (anonymous):

those are the questions, please help me!

random231 (random231):

wat is the question?

OpenStudy (anonymous):

What are you looking for. Can't solve it if I don't know.

OpenStudy (anonymous):

I am looing to solve and write each equation in slope intercept form

OpenStudy (anonymous):

looking**

OpenStudy (anonymous):

if slope intercept form is y = mx+ b, put them in that form.

random231 (random231):

yeah rearrange the equation in y=mx+c form

OpenStudy (anonymous):

I need the answer

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (anonymous):

1. -3x=-x-18

OpenStudy (anonymous):

2. 4x=-5x+100

random231 (random231):

okay ill help you with one:3x=4y+8 implies 4y=3x-8 implies y=3x/4-2

OpenStudy (anonymous):

3.-2x=-9x+13

OpenStudy (anonymous):

that is just rearranging the numbers suckerofmath

OpenStudy (anonymous):

That's all you have to do to pu it in the right form